OpenStudy (anonymous):

I'm trying to understand how ACF is actually calculated. Let's say we have the following dataset. How would we calculate the ACF of YtYt-2. t Yt 1 3 2 6 3 8 4 4 5 4 6 8

OpenStudy (amistre64):

hopefully you can better define what ACF stands for.

OpenStudy (amistre64):

\bar y \[\bar y\]

OpenStudy (anonymous):

ACF is autocorrelation function

OpenStudy (anonymous):

The ACF function is below for reference: \[\frac{\sum_{t=1+k}^{n} (Y _{t}-\bar Y)(Y _{t-k}-\bar Y) }{ \sum_{t = 1}^{n} (Y_{t} - \bar Y)^2}\]] \]

OpenStudy (amistre64):

http://mathworld.wolfram.com/Autocorrelation.html autocorrelation is beyond the scope of my ken. can you make any sense of the wolfram documentation?

OpenStudy (anonymous):

Unfortunately not regarding the Wolfram documentation. Thanks for taking a look.

OpenStudy (amistre64):

good luck with it :) there are plenty of smarter than mes about that may have a better head on their shoulders for this topic :)